A project has the following estimated data: price = $52 per unit; variable costs = $33 per unit; fixed costs = $15,500; required return = 12 percent; initial investment = $32,000; life = four years. Ignoring the effect of taxes, what is the accounting break-even quantity? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Break-even quantity What is the cash break-even quantity? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Break-even quantity What is the financial break-even quantity? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Break-even quantity What is the degree of operating leverage at the financial break-even level of output? (Do not round intermediate calculations. Round your answer to 3 decimal places, e.g., 32.161.) DOL
Answer #1

Using a straight line depreciation method, depreciation per year = ($32,000/4) = $8,000. Hence, since there are no taxes,

Accounting break-even quantity = (Fixed costs + Depreciation)/ (sales price per unit - variable costs per unit)

= ($15,500 + $8,000)/($52-$33) = $23,500/$19 = 1236.84 units

Cash break even quantity does not count depreciation which is a non-cash item.

Hence, cash break even quantity = ($15,500 + $8,000)/($52-$33) = $15,500/$19 = 815.79 units

The financial break-even point is the level at which earnings before interest and taxes or EBIT needs to be for the net profit to be $0. Since, the above mentioned question omits taxes and there is no interest mentioned, the Financial break even point is $0

EBIT is $0 at ($15,500 + $8,000)/($52-$33) = $23,500/$19 = 1236.84 units

At this point the degree of operating leverage = [1,236.84($52-$33)]/[1,236.84($52-$33)-$15,500] = 2.937